a new self-tanning lotion for everyday use is to be sold .first ,an experimental lotion mixture is made by mixing 600 onces of everyday moisturizing lotion worth $0.60 an once with self-tanning lotion worth $3 per once.if the experimental lotion is to cost $1.20 per once,how many onces of the self tanning lotion should be in the mixture?
Answer by jorel1380(3719)

600(.60)+n(3)=1.20(600+n)
360+3n=720+1.2n
1.8n=360
n=360/1.8=200 ozs. of self-tanning lotion necessary..